OpenStudy (anonymous):

Find the points on the parabola y=x2 that are closest to the point (0,5).

OpenStudy (anonymous):

ok so the length of the line from (0,5) to a point on the parabola x² becomes a function (rise/run): (5-x²)/(0-x)=5-x²/x

OpenStudy (anonymous):

that can be differentiated then to find the maximum value point on the f(x)=5-x²/x: f'(x)=-x²-5/x²

OpenStudy (jack1):

Let the required point be (p, p^2). Distance from P to (0, 5) squared D^2 = (p - 0))^2 + (p^2 - 5)^2 = (p)^2 + (p^2 - 5)^2 Multiply out and differentiate to get 2D*dD/dp = f(p) For minimum distance solve f(p) = 0.

OpenStudy (jack1):

the points i got were: \[x = (\sqrt{2} \times-3) / 2 or x = (2√×3)/2\]

OpenStudy (jack1):

9/2, (-3 x sqrt2)/2

OpenStudy (anonymous):

@Jack1 can you clarify 2D*dD/dp ? I didn't follow

OpenStudy (anonymous):

\[d^2=x^2+(x^2-5)^2\] \[d^2=x^2+x^4-10x^2+25=x^4-9x^2+25\]

OpenStudy (anonymous):

derivative is \(4x^3-18x\) which is zero is \(x=0\) or if \(4x^2-18=0\)

OpenStudy (anonymous):

second one gives \(x^2=\frac{9}{2}\)

OpenStudy (anonymous):

@satellite73 why is it ok to use the square of the distance?

OpenStudy (amistre64):

because the ratios of x y and d remain the same ...

OpenStudy (amistre64):

since d = sqrt(x^2+y^2, then d^2 = x^2 + y^2

OpenStudy (amistre64):

my idea is to take the derivative; 2x, as the general slope of a tangent line; therefore we want the perp line (for closest point) to be -1/2x for any point (a,b) on the function \[y=-\frac{1}{2a}(x-a)+b\]to make sure the line has the stated point (0,5), plug it in \[0=-\frac{1}{2a}(5-a)+b\] \[0=-\frac{5}{2a}+\frac52+b\] \[b=\frac{5}{2a}-\frac52\]set y=x^2 to a,b and equate \[a^2-b=\frac{5}{2a}-\frac52-b\]

OpenStudy (amistre64):

im sure i messed that up along the way, but it was a good effort lol

OpenStudy (amistre64):

yes, i used (0,5) as (y,x) instead of (x,y) :/

OpenStudy (amistre64):

\[5=-\frac{1}{2a}(0-a)+b\] \[5=\frac12+b\] \[b=4.5\]therefore y = 4.5 when x = sqrt(4.5) is my guess

OpenStudy (anonymous):

@amistre64 what does the solution x=0 mean when the point (0,y) is clearly not one of the nearest points?

OpenStudy (anonymous):

ah it's a local maximum?

OpenStudy (amistre64):

x=0 is a line

OpenStudy (amistre64):

x=0 is also known as the y axis, they represent the same graphical line

OpenStudy (anonymous):

I mean, if the derivative has three minimum or maximum points (-sqrt(9/2), 0, sqrt(9/2)), how do I check which ones they are? I know intuitively

OpenStudy (amistre64):

the derivative =0, or undefined, give us critical values that we need to check by plugging those critical values back into the original setup. we can also use another derivative and check for sign changes if need be.

OpenStudy (amistre64):

btw, the derivative of x^2 is just 2x, and has only one critical point, 2x = 0 when x=0

OpenStudy (amistre64):

the derivative of 2x is 2, which is always positive and tells us the slope is increasing as x approaches infinity